# Free Doubly-Infinitary Distributive Categories are Cartesian Closed

### v6

@@ 4.6. Category of topological spaces
The category $\mathbf{FinSet}$ of finite sets and functions. It does not have infinite products and coproducts, so is not doubly-infinitary distributive.
@@
More interestingly, there exist categories that have all products and coproducts, are infinitary distributive and cartesian closed, but still fail to be doubly-infinitary distributive. Consider the category $\mathbf{Qbs}$ of quasi-Borel spaces [20]. $\mathbf{Qbs}$ is the category of concrete sheaves over standard Borel spaces with countable measurable covers, hence, as a Grothendieck quasi-topos, it is complete, cocomplete and cartesian closed. It is of interest as a generalized setting for probability theory as it is a concrete category that has a full, limit and countable coproduct preserving embedding of standard Borel spaces into it.
Concretely, a quasi-Borel space $X$ consists of of a set $|X|$ and a set $M_{X}$ of maps $\mathbb{R}\to|X|$ such that (1) $M_{X}$ contains all constant functions; (2) $M_{X}$ is closed under precomposition with measurable functions $\mathbb{R}\to\mathbb{R}$; (3) $M_{X}$ is closed under gluing along countable measurable partitions of $\mathbb{R}$. Morphisms $X\to Y$ are functions $f:|X|\to|Y|$ such that $f\circ g\in M_{Y}$ if $g\in M_{X}$.
We have products $\prod_{j\in J}X_{j}$ given by $|\prod_{j\in J}X_{j}|=\prod_{j\in J}|X_{j}|$ and $M_{\prod_{j\in J}X_{j}}=\{\langle f_{j}\mid j\in J\rangle\mid f_{j}\in M_{X_{j}}\}$ and coproducts $\coprod_{j\in J}X_{j}$ given by $|\coprod_{j\in J}X_{j}|=\coprod_{j\in J}|X_{j}|$ and $M_{\coprod_{j\in J}X_{j}}=\{\lambda r.\iota_{j(n(r))}(f_{n(r)}(r))\mid n:\mathbb{R}\to\mathbb{N}\text{ measurable },j:\mathbb{N}\to J,\forall m\in\mathbb{N}.f_{m}\in M_{X_{j(m)}}\}$. Observe that elements of $M_{\coprod_{j\in J}X_{j}}$ always factor over countably many components of the coproduct. Countable coproducts are well-behaved in the sense that $M_{\coprod_{j\in J}X_{j}}=\{\lambda r.\iota_{j(r)}(f_{j(r)}(r))\mid j:\mathbb{R}\to J\text{ measurable },f_{j}\in M_{X_{j}}\}$ if $J$ is countable.
The question is whether $d:\left(\prod_{j\in J}\coprod_{i\in I_{j}}C_{ij}\right)\to\left(\coprod_{f:\prod_{j\in J}I_{j}}\prod_{j\in J}C_{f(j)j}\right)$ is a morphism of $\mathbf{Qbs}$. We need to show that $d\circ g\in M_{\coprod_{f:\prod_{j\in J}I_{j}}\prod_{j\in J}C_{f(j)j}}$ for all $g\in M_{\prod_{j\in J}\coprod_{i\in I_{j}}C_{ij}}$. This is clearly true iff $\{j\in J\mid\#I_{j}<2\}$ is finite. In particular, $\mathbf{Qbs}$ is infinitary distributive, but not doubly-infinitary distributive.
@@ 5. Final remarks and future work
We discussed the concept of doubly-infnintary distributive categories, which is a a notion that naturally arises from a canonical pseudodistributive law between the free product completion pseudomonad and the free coproduct completion pseudomonad. The most surprising result is that free doubly-infinitary distributive category $\mathbf{Dist}(\mathbb{C})$ on a category $\mathbb{C}$ is cartesian closed.
We could also consider free completions under finite coproducts and products. This would amount to replacing $\mathbf{Fam}$ with $\mathbf{FinFam}$, where $\mathbf{FinFam}(\mathbb{C})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\displaystyle\Sigma_{\mathbf{FinSet}}\mathfrak{Fam}(\mathbb{C})$. The distributive law
$\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C})^{op})^{op}\to\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})$
then gives rise to the theory of distributive categories in the usual sense (i.e., categories with finite products and coproducts that distribute), while the distributive law
$\mathbf{FinFam}(\mathbf{Fam}(\mathbb{C})^{op})^{op}\to\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{op})^{op})$
gives rise to the theory of infinitary distributive categories in the usual sense (i.e., categories with finite products and infinite coproducts that distribute. As a small variation on our results, we get that $\mathbf{FinDist}(()\mathbb{C})=\mathbf{FinFam}((\mathbf{FinFam}(\mathbb{C}^{op}))^{op})$ is cartesian closed for any locally finite category $\mathbb{C}$.
@@ 5.1. Further properties of $\mathbf{Dist}(\mathbb{C})$
We remark that $\mathbf{Dist}(\mathbb{C})$ has many other interesting properties, including most notably extensivity. Many of these properties are already in the literature, since they follow from the fact that $\mathbf{Dist}{\mathbb{C}}=\mathbf{Fam}{\mathbb{D}}$ is the free coproduct completion of $\mathbb{D}=\mathbf{Fam}{\mathbb{C}^{\mathrm{op}}}^{\mathrm{op}}$. We refer to [10, 8, 3, 33] for some of these properties.
@@ 5.2. Non-canonical isomorphisms
In Definition 3.1, doubly-infinitary distributive categories are defined in terms of a canonical morphism being invertible. It is natural to ask, then, if finding a non-canonical isomorphism would suffice (this is Pisani’s question for distributive categories, e.g. [25]).
@@
$\left(\coprod_{f\in\Pi_{j\in J}I_{j}}\prod_{j\in J}C_{f_{j}j}\right)\to\left(\prod_{j\in J}\coprod_{i\in I_{j}}C_{ij}\right).$
for each family of objects $\left(\left(C_{ij}\right)_{i\in I_{j}}\right)_{j\in J}$.
Finally, since a coproduct and product complete category is doubly-ifninitary distributive if and only if the coproduct functor $\coprod:\mathbf{Fam}{\mathbb{C}}\to\mathbb{C}$ is product-preserving, we can achieve the non-canonical result stated above by results on preservation of limits from naturality (non-canonical isomorphisms), see, for instance, [31, 9].
@@ 5.3. Generalized categorical structures
It was recently established that the category of $\left(T,\mathbb{V}\right)$-categories is extensive, see [11]. This shows that a plethora of categories of topological categories and, more generally, (enriched) categorical structures share the property of being extensive.
In exploring the realm of doubly-infinitary distributivity, we unveiled that both the category of locally connected spaces and the category of (small) categories possess this property. However, notably, the category of topological spaces does not share this trait. This observation leads us to pose the following open question:
Open question 5.1.
It remains an open question: under which conditions on $T$ and $\mathbb{V}$ does the category of $(T,\mathbb{V})$-categories exhibit doubly-infinitary distributivity?
@@
By 4.13, the obvious road map to solve the problem above would start with a characterization of the categories $(T,\mathbb{V})$-$\mathbf{Cat}$ of $(T,\mathbb{V})$-categories that are free doubly-infinitary distributive on its category of connected objects. Enlightened by [12, Corollary 4.7], this happens to be the case under suitable conditions, for $\mathbb{V}$ thin and intersection-preserving $T$.
@@ 5.4. Gödel’s Dialectica interpretation
The exponential formula for $\mathbf{Dist}(\mathbb{C})$ presented in this paper is a variation on Gödel’s Dialectica interpretation (see [18] for the original reference and [21] for a categorical formulation), exponentials in categories of polynomials/containers [5], and the Diller-Nahm formulas (see [16] for the original reference and [21] and [33, Section 6.4] for categorical formulations). We explain the precise relationship to these results in the separate paper [26], in order not to distract from the simple nature of the results presented herein.

### v7

@@ 5.7. Cartesian closed categories are not necessarily doubly-infinitary distributive
Despite being cartesian closed, the category $\mathbf{FinSet}$ of finite sets and functions. It does not have infinite products and coproducts, so is not doubly-infinitary distributive.
@@
More interestingly, there exist categories that have all products and coproducts, are infinitary distributive and cartesian closed, but still fail to be doubly-infinitary distributive. Consider the category $\mathbf{Qbs}$ of quasi-Borel spaces [22].
$\mathbf{Qbs}$ is a complete, cocomplete quasitopos and, hence, cartesian closed. It is of interest as a generalized setting for probability theory as it is a concrete category that has a full, limit and countable coproduct preserving embedding of standard Borel spaces into it.
The category of standard Borel spaces is not doubly-infinitary distributive, and therefore neither is $\mathbf{Qbs}$. Therefore $\mathbf{Qbs}$ is an example of a cartesian closed category that is not doubly-infinitary distributive.
Remark 5.14.
For the interested reader, we spell out the details concerning quasi-Borel spaces. Concretely, a quasi-Borel space $X$ consists of a set $|X|$ together with a set $M_{X}$ of maps $\mathbb{R}\to|X|$ satisfying the following conditions:
  1. (1)
    $M_{X}$ contains all constant functions;
  2. (2)
    $M_{X}$ is closed under precomposition with measurable functions $\mathbb{R}\to\mathbb{R}$;
  3. (3)
    $M_{X}$ is closed under gluing along countable measurable partitions of $\mathbb{R}$.
  • Morphisms $X\to Y$ are functions $f:|X|\to|Y|$ such that $f\circ g\in M_{Y}$ whenever $g\in M_{X}$.
Products $\prod_{j\in J}X_{j}$ are given by
$|\textstyle\prod_{j\in J}X_{j}|=\prod_{j\in J}|X_{j}|,\quad M_{\prod_{j\in J}X_{j}}=\{\langle f_{j}|j\in J\rangle\mid f_{j}\in M_{X_{j}}\}.$
Coproducts $\coprod_{j\in J}X_{j}$ are given by
$|\textstyle\coprod_{j\in J}X_{j}|=\coprod_{j\in J}|X_{j}|,\quad M_{\coprod_{j\in J}X_{j}}=\{\,\lambda r.\,\iota_{j(n(r))}(f_{n(r)}(r))\mid n:\mathbb{R}\to\mathbb{N}\text{ measurable},\;j:\mathbb{N}\to J,\;\forall m\in\mathbb{N}.\,f_{m}\in M_{X_{j(m)}}\,\}.$
Observe that elements of $M_{\coprod_{j\in J}X_{j}}$ always factor through countably many components of the coproduct. Countable coproducts are well behaved in the sense that
$M_{\coprod_{j\in J}X_{j}}=\{\,\lambda r.\,\iota_{j(r)}(f_{j(r)}(r))\mid j:\mathbb{R}\to J\text{ measurable},\;f_{j}\in M_{X_{j}}\,\}$
whenever $J$ is countable.
We now ask whether
$d:\Bigl(\prod_{j\in J}\coprod_{i\in I_{j}}C_{ji}\Bigr)\to\Bigl(\coprod_{f:\prod_{j\in J}I_{j}}\prod_{j\in J}C_{jf(j)}\Bigr)$
is a morphism in $\mathbf{Qbs}$. That is, we must show that $d\circ g\in M_{\coprod_{f:\prod_{j\in J}I_{j}}\prod_{j\in J}C_{jf(j)}}$ for every $g\in M_{\prod_{j\in J}\coprod_{i\in I_{j}}C_{ji}}$. This holds if and only if the set $\{\,j\in J\mid\#I_{j}\geq 2\,\}$ is finite.
This shows that $\mathbf{Qbs}$ is infinitary distributive but not doubly-infinitary distributive.
Counterexample 5.15 (Pseudotopological spaces).
Let $\mathbf{PsTop}$ denote the category of pseudotopological spaces (see, for example, [23]), which is a cocomplete and complete quasitopos, so cartesian closed.
Recall that the inclusion of $\mathbf{Top}$ into $\mathbf{PsTop}$ preserves products and coproducts, hence it follows immediately that $\mathbf{PsTop}$ is not doubly-infinitary distributive.
Therefore, $\mathbf{PsTop}$ is a cartesian closed category that is not doubly-infinitary distributive.
@@ 6. Final remarks and future work
We discussed the concept of doubly-infinitary distributive categories, a notion that naturally arises from a canonical pseudodistributive law between the free product completion pseudomonad and the free coproduct completion pseudomonad. Recall that considering free completions under finite coproducts and products amounts to replacing $\mathbf{Fam}$ with $\mathbf{FinFam}$, where $\mathbf{FinFam}(\mathbb{C})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Sigma_{\mathbf{FinSet}}\mathfrak{Fam}(\mathbb{C})$.
The distributive law
$\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C})^{\mathrm{op}})^{\mathrm{op}}\to\mathbf{FinFam}(\mathbf{FinFam}(\mathbb{C}^{\mathrm{op}})^{\mathrm{op}})$
then gives rise to the theory of distributive categories in the usual sense (that is, categories with finite products and coproducts that distribute), while the distributive law
$\mathbf{FinFam}(\mathbf{Fam}(\mathbb{C})^{\mathrm{op}})^{\mathrm{op}}\to\mathbf{Fam}(\mathbf{FinFam}(\mathbb{C}^{\mathrm{op}})^{\mathrm{op}})$
gives rise to the theory of infinitary distributive categories in the usual sense (that is, categories with finite products and infinite coproducts that distribute).
Recall that free doubly-infinitary distributive category $\mathbf{Dist}(\mathbb{C})$ on a category $\mathbb{C}$ is cartesian closed by Theorem 2.4. As a small variation on our results, we obtain that $\mathbf{FinDist}(\mathbb{C})=\mathbf{FinFam}((\mathbf{FinFam}(\mathbb{C}^{\mathrm{op}}))^{\mathrm{op}})$ is cartesian closed for any locally finite category $\mathbb{C}$.
@@ 6.1. Further properties of $\mathbf{Dist}(\mathbb{C})$
We remark that $\mathbf{Dist}(\mathbb{C})$ possesses many other interesting properties, most notably extensivity. Many of these properties are already established in the literature, as they follow from the fact that $\mathbf{Dist}(\mathbb{C})=\mathbf{Fam}(\mathbb{D})$ is the free coproduct completion of $\mathbb{D}=\mathbf{Fam}(\mathbb{C}^{op})^{op}$. Indeed, the literature on properties of the free coproduct completion of a category is quite extensive; we refer the interested reader to [3, 8, 10, 36, 37, 44, 46, 48] for discussions of some of these properties.
@@ 6.2. Non-canonical isomorphisms
Definition 3.1 introduces doubly-infinitary distributive categories in terms of the invertibility of a canonical morphism. It is then natural to ask whether the existence of a non-canonical isomorphism would suffice (this is Pisani’s question for distributive categories; see, for example, [28]).
@@
$\left(\coprod_{f\in\prod_{j\in J}I_{j}}\prod_{j\in J}C_{f_{j}j}\right)\to\left(\prod_{j\in J}\coprod_{i\in I_{j}}C_{ji}\right).$
for each family of objects $\left(\left(C_{ji}\right)_{i\in I_{j}}\right)_{j\in J}$.
Finally, since a coproduct and product complete category is doubly-infinitary distributive if and only if the coproduct functor $\coprod:\mathbf{Fam}{\mathbb{C}}\to\mathbb{C}$ is product-preserving, we can achieve the non-canonical result stated above by results on preservation of limits from naturality (non-canonical isomorphisms), see, for instance, [9, 34].
@@ 6.3. Generalized categorical structures
It was recently established that the category of $\left(T,\mathbb{V}\right)$-categories is extensive; see [12]. This result shows that a wide range of categories of topological categories and, more generally, (enriched) categorical structures share the property of being extensive.
In this work, we have shown that both the category of locally connected topological spaces and the category of (small) categories are doubly-infinitary distributive. However, notably, the category of topological spaces itself does not share this property. This observation naturally leads us to pose the following open question:
Open Question 6.1.
Under which conditions on $T$ and $\mathbb{V}$ does the category of $(T,\mathbb{V})$-categories exhibit doubly-infinitary distributivity?
@@
By 5.10, the obvious road map to solve the problem above would start with a characterization of the categories $(T,\mathbb{V})$-$\mathbf{Cat}$ of $(T,\mathbb{V})$-categories that are free doubly-infinitary distributive on its category of connected objects. Enlightened by [13, Corollary 4.7], this happens to be the case under suitable conditions, for $\mathbb{V}$ thin and intersection-preserving $T$.
@@ 6.4. Gödel’s Dialectica interpretation
The exponential formula for $\mathbf{Dist}(\mathbb{C})$ presented in this paper is a variation on Gödel’s Dialectica interpretation (see [20, 24, 17, 52]), exponentials in categories of polynomials/containers [5], and the Diller-Nahm formulas (see [18] for the original reference and [24] and [37, Section 6.4] for categorical formulations). We explain the precise relationship to these results in the separate paper [29], in order not to distract from the simple nature of the results presented herein.